Hankins Corporation has 5.7 million shares of common stock outstanding, 306,000 shares of 4.3 percent preferred stock outstanWhat is the firms cost of each form of financing? (Do not round intermediate calculations and enter your answers as a percenthat is the one question, solve all of them please
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=73.2*5700000
=417240000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*165000*1.04
=171600000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=104.6*306000
=32007600
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=417240000+171600000+32007600
=620847600
Weight of equity = MV of Equity/MV of firm
Weight of equity = 417240000/620847600
W(E)=0.672
Weight of debt = MV of Bond/MV of firm
Weight of debt = 171600000/620847600
W(D)=0.2764
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 32007600/620847600
W(PE)=0.0516
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 3.5 + 1.13 * (6.9)
Cost of equity% = 11.3
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =22x2
1040 =∑ [(5.3*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^22x2
                   k=1
YTM = 4.9982061648
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 4.9982061648*(1-0.23)
= 3.848618746896 = 3.85%
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 4.3/(104.6)*100
=4.11
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.85*0.2764+11.3*0.672+4.11*0.0516
WACC =8.87%